OpenStudy (anonymous):

evaluate:∫v dv/(6v²-1) again plz help me to solve this problem,,tnx

OpenStudy (lalaly):

let u = 6v^2 -1 du=12vdv substituting 1/12∫du/u = 1/12 lnu substitute v again 1/12 ln(6v^2 -1)

OpenStudy (anonymous):

tnx

OpenStudy (lalaly):

:)